### v1

@@
A Strategic Model of Software Dependency Networksthanks: This version: February 20, 2024. Contact: cfritz@psu.edu, co-pierre.georg@uct.ac.za, angelo.mele@jhu.edu, and michael.schweinberger@psu.edu. We wish to thank participants at the Network Science in Economics 2023 at Virginia Tech, 2023 BSE Summer Forum Networks Workshop, Southern Economic Association Annual Meeting 2023 as well as seminar participants at Georgetown University and UPenn for helpful comments and suggestions. The authors acknowledge support from the German Research Foundation award DFG FR 4768/1-1 (CF), Ripple’s University Blockchain Research Initiative (CPG), the U.S. Department of Defense award ARO W911NF-21-1-0335 (MS, CF), and the U.S. National Science Foundation awards NSF DMS-1513644 and NSF DMS-1812119 (MS). Mele acknowledges support from an IDIES Seed Grant from the Institute for Data Intensive Engineering and Science at Johns Hopkins University and partial support from NSF grant SES 1951005.
Cornelius Fritz The Pennsylvania State University – Department of Statistics Co-Pierre Georg University of Cape Town Angelo Mele Johns Hopkins University – Carey Business School Michael Schweinberger The Pennsylvania State University – Department of Statistics
Modern software development involves collaborative efforts and reuse of existing code, which reduces the cost of developing new software. However, reusing code from existing packages exposes coders to vulnerabilities in these dependencies. We study the formation of dependency networks among software packages and libraries, guided by a structural model of network formation with observable and unobservable heterogeneity. We estimate costs, benefits, and link externalities of the network of $696,790$ directed dependencies between $35,473$ repositories of the Rust programming language using a novel scalable algorithm. We find evidence of a positive externality exerted on other coders when coders create dependencies. Furthermore, we show that coders are likely to link to more popular packages of the same software type but less popular packages of other types. We adopt models for the spread of infectious diseases to measure a package’s systemicness as the number of downstream packages a vulnerability would affect. Systemicness is highly skewed with the most systemic repository affecting almost 90% of all repositories only two steps away. Lastly, we show that protecting only the ten most important repositories reduces vulnerability contagion by nearly 40%.
Keywords: software development, dependency graphs, strategic network formation, exponential random graphs
JEL Classification: D85, L17, L86, O31
@@ 1 Introduction
Modern software development is a collaborative effort that results in sophisticated software packages that makes extensive use of already existing software packages and libraries (Schueller et al., 2022). This results in a complex network of dependencies among software packages, best described as dependency graphs. [^1] Dependency graphs have been studied for many different programming languages using information from package managers. For an empirical comparison of the most common ones, see Decan et al. (2019). While this results in significant efficiency gains for software developers (Lewis et al., 1991; Barros-Justo et al., 2018), it also increases the risk that a vulnerability in one software package renders a large number of other software packages also vulnerable. In line with this view, the number of vulnerabilities listed in the Common Vulnerabilities and Exposure Database has increased almost twelve-fold between 2001 and 2019.222The Common Vulnerabilities and Exposure database is a public reference for known information-security vulnerabilities which feeds into the U.S. National Vulnerability Database. See https://cve.mitre.org/. The reported number of incidents has increased from 3,591 for the three years from 1999 to 2001 to 43,444 for the three-year period from 2017 to 2019. A recent estimate by Krasner (2018) put the cost of losses from software failures at over USD 1 Trillion, up from USD 59 Billion in 2002, estimated by the National Institute of Standards and Technology (2002). [^3] Industry group Cybersecurity Ventures estimates that the damages incurred by all forms of cyber crime, including the cost of recovery and remediation, totalled USD 6 Trillion in 2021, and could reach USD 10.5 Trillion annually by 2025 (Source).
@@
Because vulnerabilities can spread from one software package to all packages depending on it, coders can create an externality when deciding to re-use code from existing packages rather than implementing the required functionality themselves. Whether this is the case and how large the externality is, is ultimately an empirical question. We model coders’ decisions to create a dependency on another software package as an equilibrium game of network formation with observed and unobserved heterogeneity. The equilibrium characterization provides a likelihood of observing a particular architecture of the software dependency network in the long run. Using data on open source software projects of the Rust programming language, we obtain all $696,790$ dependencies among $35,473$ repositories of software projects contained in Rust’s software package manager Cargo. We find evidence that a maintainer’s decision to allow a dependency exerts a positive externality on other maintainers: Highly interdependent software packages are likely to become even more interdependent. This means that it is particularly important to ensure that such packages are free of vulnerabilities that potentially can affect a large number of other software packages.
@@
Open source software is an ideal laboratory to study software dependencies. Since the actual source code of a software package can be inspected and modified by others, code reuse is not only possible, it is encouraged. Modern software packages are not developed in isolation, but within an ecosystem of interdependent packages. A package manager keeps track of code dependencies and ensures that these are automatically resolved so that when a user installs one package, all of the package’s dependencies are also installed and automatically satisfied. There are many programming languages which provide software using package managers. We restrict our analysis to repositories of software packages written in the programming language Rust and managed by the Cargo package manager. [^4] Cargo is Rust’s official package manager. See here for the official documentation. Rust is a modern programming language that allows particularly safe software development without compromising performance. [^5] Rust has been voted the most loved programming language by developers on the popular coding knowledge exchange website http://stackoverflow.com. We use data from https://libraries.io, which, for each software repository includes a full list of all dependencies.
@@
We model the development of software as a process where coders reuse existing code, creating a network of dependencies among software packages—organized in repositories, which can include several closely related software packages—in the process. This is typical for today’s prevalent object-oriented software development paradigm. Our model describes a system of $N$ individual software repositories. Each repository is managed by a group of coders who decide which dependencies to form. Coders obtain a net benefit of linking to another package which depends on how active this dependency is maintained, how mature, popular, and large it is. We also allow a coders’ utility to be affected by a local, i.e. type-specific, externality: since dependent packages have dependencies themselves, they are susceptible to vulnerabilities imported from other packages. So we assume that coders care about the direct dependencies of the packages they link to. This specification excludes externalities that are more than two links away as in the network formation models of DePaula et al. (2018), Mele (2017) and Mele and Zhu (forthcoming). The network of dependencies forms over time and in each period a randomly selected package needs an update, so the coders decide whether to form a link or update the software in-house. Before updating the link, the package receives a random match quality shock. We characterize the equilibrium distribution over networks as a mixture of exponential random graphs (Schweinberger and Handcock, 2015; Mele, 2022), which can be decomposed into within- and between-type contribution to the likelihood.
@@
Estimation of this model is complicated because the likelihood depends on a normalizing constant that is infeasible to compute in large networks. Moreover, the model’s unobserved heterogeneity has to be integrated out in the likelihood, thus further complicating the computations. To alleviate these problems, we resort to an approximate two-step approach. In the first step, we estimate the discrete types of the nodes through approximations of the likelihood via a variational mean-field algorithm for stochastic blockmodels (Vu et al., 2013). In the second step, we estimate the structural payoff parameters using a fast Maximum Pseudo-Likelihood Estimator (MPLE), conditioning on the estimated types (Babkin et al., 2020; Dahbura et al., 2021). [^6] These methods are implemented in a highly scalable open source R package bigergm, available on Github at: https://github.com/sansan-inc/bigergm. See the Appendix for details on the implementation.
@@
Our main result is that coders exert a positive externality on other coders when creating a dependency on another package. Other coders are then also more likely to create a dependency with that package. As a result, packages that are already highly interdependent tend to become even more interdependent. This increases the risk that a vulnerability in a single package has large adverse consequences for the entire ecosystem. One example of how a vulnerability in one package impacted a significant part of critical web infrastructure is Heartbleed, the infamous bug in the widely used SSL/TLS cryptography library, resulting from an improper input variable validation (Durumeric et al., 2014). When Heartbleed was disclosed, it rendered up to 55% of secure webservers vulnerable to data theft, including of the Canadian Revenue Agency, and Community Health Systems, a large U.S. hospital chain. Our model provides a further argument for ensuring the security of highly interdependent software packages. Not only can a vulnerability in such a package affect a large fraction of the entire ecosystem, because of the externality we identify, it is also likely that this fraction increases as time goes on.
@@
Another important question in the study of network formation processes is whether linked packages are similar or dissimilar with respect to their covariates. We find evidence of similarity among the same type of software packages in terms of their Size, Popularity, and Maturityas well as between different-type software packages in terms of their Size. We further find evidence of dis-similarity between different-type software packages in terms of their Popularity, Num.Contributors (as an alternative popularity measure), and Num.Forks. In other words, mature, large and popular software packages of one type are likely to depend on similar packages of the same type, but on less popular and mature software packages of another type. This is intuitive, because mature, popular, and large packages are likely to have a large user base with high expectations of the software and coders try to satisfy this demand by providing sophisticated functionality with the help of mature, popular, and large software packages of the same type. These packages are often built using a lot of low-level functionality from large but fairly generic libraries, which is why large popular and mature software packages of one type depend on larger, but less popular and less mature software packages of another type. One example of this is the inclusion of a payment gateway in an e-commerce application. The e-commerce application itself can be sophisticated and complex, like ebay is in the provision of their auction mechanism. For such an application it is particularly valuable to provide users with the ability to use a variety of different payment methods, including credit card, paypal, or buy now pay later solutions like Klarna. Our result is also in line with the recent trend in software development away from large monolithic applications and towards interconnected microservices (Traore et al., 2022).
Lastly, we adapt the susceptible-infected-recovered (SIR) model from epidemiology to examine the spread of vulnerabilities in the Rust Cargo dependency graph. We measure a repository’s $k$-step systemicness as the number of downstream packages that are potentially rendered vulnerable by a vulnerability in the upstream repository. We find that systemicness is highly skewed across repositories: While the average vulnerable repository affects $52$ repositories two steps away, the most systemic repository affects $26,631$ repositories. This heterogeneity underscores the existence of key nodes within the software dependency graph whose compromise could lead to widespread vulnerability exposure. We then study the efficacy of targeted interventions to mitigate vulnerability spread, drawing parallels with vaccination strategies in public health. By focusing on securing the most critical nodes—determined by a combination of betweenness centrality and expected fatality—we assess the potential to significantly reduce the risk of vulnerability contagion. Our findings suggest that protecting the top ten most critical repositories can reduce the extent of vulnerability contagion by nearly 40%.
@@
The package manager model is nowadays adopted by most programming languages which makes it feasible to use dependency graphs to study a wide variety of settings. Kikas et al. (2017), for example, study the structure and evolution of the dependency graph of JavaScript, Ruby, and also Rust. The authors emphasize that dependency graphs of all three programming languages have become increasingly vulnerable to the removal of a single software package. An active literature studies the network structure of dependency graphs (Decan et al., 2019) to assess the vulnerability of a software ecosystem (see, for example, Zimmermann et al. (2019)).
@@
These papers show the breadth of literature studying open source ecosystems and dependency graphs. However, the literature considers the network either as stochastic or even as static and given. By contrast, we model the formation of dependencies as coders’ strategic choice in the presence of various and competing mechanisms that either increase or reduce utility. [^9] Blume et al. (2013) study how possibly contagious links affect network formation in a general setting. While we do not study the consequences of the externality we identify for contagion, this is a most worthwhile avenue for future research.
@@
The theoretical literature on strategic network formation has pointed out the role of externalities in shaping the equilibrium networks (Jackson, 2008; Jackson and Wolinsky, 1996). Estimating strategic models of network formation is a challenging econometric task, because the presence of externalities implies strong correlations among links and multiple equilibria (Mele, 2017; Snijders, 2002; DePaula et al., 2018; Chandrasekhar, 2016; DePaula, 2017; Boucher and Mourifie, 2017; Graham, 2017, 2020). In this paper, we model network formation as a sequential process and focus on the long-run stationary equilibrium of the model (Mele, 2017, 2022). Because the sequential network formation works as an equilibrium selection mechanism, we are able to alleviate the problems arising from multiple equilibria.
@@
Adding unobserved heterogeneity further complicates identification, estimation and inference (Schweinberger and Handcock, 2015; Graham, 2017; Mele, 2022). Other works have considered conditionally independent links without externalities (Graham, 2017, 2020; DePaula, 2017; Chandrasekhar, 2016), providing a framework for estimation and identification in random and fixed effects approaches. On the other hand, because link externalities are an important feature in this contexts, we move away from conditionally independent links, and model nodes’ unobserved heterogeneity as discrete types, whose realization is independent of observable characteristics and network, in a random effect approach. We can thus adapt methods of community discovery for random graphs to estimate the types, following approximations of the likelihood suggested in Babkin et al. (2020) and improved in Dahbura et al. (2021) to accomodate for observed covariates. Our two-steps method scales well to large networks, thus improving the computational challenges arising in estimation of these complex models (Boucher and Mourifie, 2017; Vu et al., 2013; Bonhomme et al., 2019).
@@
Our model is able to identify externalities as well as homophily (Currarini et al., 2010; Jackson, 2008; Chandrasekhar, 2016), the tendency of individuals to form links to similar individuals. On the other hand, our model can also detect heterophily (or competition) among maintainers. We also allow the homophily to vary by unobservables, while in most models the homophily is estimated only for observable characteristics (Currarini et al., 2010; Schweinberger and Handcock, 2015; DePaula et al., 2018; Chandrasekhar, 2016; Graham, 2020).

### v2

@@
Vulnerability Webs:
Systemic Risk in Software Networksthanks: The order of authors is alphabetical. fritzc@tcd.ie, co-pierre.georg@uct.ac.za, angelo.mele@jhu.edu, and mus47@psu.edu. This paper supersedes a previous version titled "A Strategic Model of Software Dependency Networks". We wish to thank Bryan Graham, Stephane Bonhomme, Louise Laage, Matt Jackson, Eric Auerbach for helpful suggestions. We also thank participants at 2024 Cowles Foundation Conference in Econometrics, ESIF 2024 Conference on ML+AI, Empirical Methods in Game Theory Workshop, MEG 2024, EC24, MGOW 2024, Network Science in Economics 2023, 2023 BSE Summer Forum Networks Workshop, SEA 2023, Econometrics of Peer Effects and Networks, and seminar participants at Georgetown University, University of Washington, Virginia Tech, Boise State University, and the University of Pennsylvania for helpful comments and suggestions. The authors acknowledge support from the German Research Foundation award DFG FR 4768/1-1 (CF), Ripple’s University Blockchain Research Initiative (CPG), the U.S. Department of Defense award ARO W911NF-21-1-0335 (MS, CF), and the U.S. National Science Foundation awards NSF DMS-1513644 and NSF DMS-1812119 (MS). Mele acknowledges support from an IDIES Seed Grant from the Institute for Data Intensive Engineering and Science at Johns Hopkins University and partial support from NSF grant SES 1951005.
Cornelius Fritz Trinity College Dublin – School of Computer Science and Statistics Co-Pierre Georg Frankfurt School of Finance and Management Angelo Mele Johns Hopkins University – Carey Business School Michael Schweinberger The Pennsylvania State University – Department of Statistics
Modern software development is a collaborative effort that re-uses existing code to reduce development and maintenance costs. This practice exposes software to vulnerabilities in the form of undetected bugs in direct and indirect dependencies, as demonstrated by the Crowdstrike and HeartBleed bugs. The economic costs resulting from such vulnerabilities can be staggering. We study a directed network of 52,897 software dependencies across 16,102 Python repositories, guided by a strategic model of network formation that incorporates both observable and unobservable heterogeneity. Using a scalable variational approximation of the conditional distribution of unobserved heterogeneity, we show that outsourcing code to other software packages by creating dependencies generates negative externalities. Modeling the propagation of risk in networks of software packages as an epidemiological process, we show that increasing protection of dependencies based on popular heuristics is ineffective at reducing systemic risk. By contrast, AI-assisted coding enables developers to replace dependencies with in-house code and reduces systemic risk.
Keywords: economics of software development, dependency graphs, strategic network formation, exponential random graphs
JEL Classification: D85, L17, C31, L86, O31
@@ 1 Introduction
On July 19, 2024 a faulty update to the software Crowdstrike—a security software adopted by many companies—created worldwide panic and disruptions, paralyzing domestic and international air traffic for hours, and disrupting operations at banks, hospitals, and hotel chains. [^1] See https://www.nytimes.com/2024/07/19/business/dealbook/tech-outage-crowdstrike-microsoft.html All told, the update affected 8.5 million Windows machines worldwide, each of which had to be manually rebooted to patch up the system and restart operations. The faulty update of the software Crowdstrike was not the first instance highlighting systemic risk in software systems, and it will not be the last: e.g., in 2014 the HeartBleed bug—a vulnerability of the OpenSSL authentication library used by many websites—compromised the data of millions of users of websites such as Eventbrite, OkCupid, in addition to disrupting operations at the FBI and hospitals. Such disruptions can generate staggering economic costs: Heartbleed, for example, resulted in at least \$currency-dollar\\$\$500 million in damages. According to the risk management company Parametrix, the disruption caused by Crowdstrike cost Fortune 500 companies \$currency-dollar\\$\$5.4 billion. [^2] See Parametrix (2024).
@@
The risk of these disruptions is deeply embedded in the structure of software systems. Modern software development is a collaborative effort that produces sophisticated software, making extensive use of already existing software packages and libraries (Schueller et al., 2022). This results in a complex network of dependencies among software packages, best described as dependency graphs. [^3] Dependency graphs have been studied for many different programming languages using information from package managers. For an empirical comparison of the most common ones, see Decan et al. (2019). While this re-use of code produces significant efficiency gains for software developers (Lewis et al., 1991; Barros-Justo et al., 2018), it also increases the risk that a vulnerability in one software package renders a large number of other packages equally vulnerable. In line with this view, the number of vulnerabilities listed in the Common Vulnerabilities and Exposure Database has increased almost twelve-fold between 2001 and 2019.444The Common Vulnerabilities and Exposure database is a public reference for known information-security vulnerabilities which feeds into the U.S. National Vulnerability Database. See https://cve.mitre.org/. The reported number of incidents has increased from 3,591 for the three years from 1999 to 2001 to 43,444 for the three-year period from 2017 to 2019. A recent estimate by Krasner (2018) put the cost of losses from software failures at over \$1 trillion, up from \$59 billion in 2002 estimated by the National Institute of Standards and Technology (2002). [^5] Industry group Cybersecurity Ventures estimates that the damages incurred by all forms of cyber crime, including the cost of recovery and remediation, totalled USD 6 Trillion in 2021, and could reach USD 10.5 Trillion annually by 2025 (Source).
Because vulnerabilities can potentially spread from one software package to all packages depending on it, coders can create an externality when deciding to re-use code from existing packages rather than implementing the required functionality themselves. Whether this is the case and how large the externality is, is ultimately an empirical question.
@@
We model coders’ decisions to create a dependency to another software package as an equilibrium game of network formation with observed and unobserved heterogeneity. Developers weigh costs and benefits of forming links and revise these links at random times, when there is a need for an update. Our equilibrium characterization provides the likelihood of observing a particular architecture of the software dependency network in the long run. Using data on open source software projects contained in Python’s software package manager Pypi, we obtain $52,897$ dependencies among $16,102$ repositories. We find evidence that a maintainer’s decision to allow a dependency exerts a negative externality on other maintainers. In principle, the presence of this externality discourages the formation of dependencies and partially contributes to the sparsity of the network. Our result implies that it is particularly important to ensure that such packages are free of vulnerabilities that potentially can affect a large number of other software packages.
We can use our model to study the effect of interventions that reduce vulnerabilities of libraries that are likely to affect a large subset of the software ecosystem. We assume that vulnerabilities, coding errors or bugs spread according to an epidemic susceptible-infected-recovered (SIR) model. Our simulations show that strategies to decrease systemic risk of contagion based on interventions that target nodes according to measures of centrality or expected risk of spread, are not very effective in mitigating the average systemic risk. On the other hand we observe that the introduction of AI-assisted coding may decrease the incentives to form dependencies through a decrease in costs of producing code. This process of decreasing the density of the network will mechanically decrease the average systemic risk. We show that even moderate productivity gains from AI translate into a significant flattening of the risk profile for the software network.
@@
Open source software is an ideal laboratory to study software dependencies. Since the actual source code of a software package can be inspected and modified by others, code reuse is not only possible, it is encouraged. Modern software packages are not developed in isolation, but within an ecosystem of interdependent packages. A package manager keeps track of code dependencies and ensures that these are automatically resolved so that when a user installs one package, all of the package’s dependencies are also installed and automatically satisfied. There are many programming languages that provide software using package managers. We restrict our analysis to repositories of software packages written in the popular programming language Python and managed by the Pypi package manager.666 PyPI stands for Python Package Index and is Python’s official third-party software repository. See here for the official documentation. Python is a modern programming language that is very popular in the software development community. We use data from libraries.io, which, for each software repository includes a full list of all dependencies.
@@
We model the development of software as a process where coders reuse existing code, creating a network of dependencies among software packages—organized in repositories, which can include several closely related software packages—in the process. This is typical for today’s prevalent object-oriented software development paradigm. Our model describes a system of $N$ individual software repositories, characterized by observable and unobservable types. Each repository is managed by a group of coders who decide which dependencies to form. Coders obtain a net benefit of linking to another package which depends on how active this dependency is maintained, how mature, popular, and large it is. We also allow a coders’ utility to be affected by a local, i.e. type-specific, externality: since dependent packages have dependencies themselves, they are susceptible to vulnerabilities imported from other packages. So we assume that coders care about the direct dependencies of the packages they link to. This specification excludes externalities that are more than two links away as in the network formation models of DePaula et al. (2018), Mele (2017) and Mele and Zhu (2023). The network of dependencies forms over time and in each period a randomly selected package needs an update, so the coders decide whether to form a link or update the software in-house. Before updating the link, the package receives a random match quality shock. We characterize the equilibrium distribution over networks as a mixture of exponential random graphs (Schweinberger and Handcock, 2015; Mele, 2022), which can be decomposed into within- and between-types contribution to the likelihood.
@@
Model estimation is challenging because the likelihood depends on a normalizing constant that is infeasible to compute in large networks. Moreover, the model’s unobserved heterogeneity has to be integrated out in the likelihood, making direct computations infeasible even for moderately sized networks. To alleviate these problems, we resort to a novel approximate two-step approach. In the first step, we estimate the discrete unobservable types of the nodes by approximating their conditional distribution via a fast variational mean-field approximation first proposed for stochastic blockmodels (Vu et al., 2013). In the second step, we estimate the structural payoff parameters using a fast Maximum Pseudo-Likelihood Estimator (MPLE), conditioning on the estimated types (Babkin et al., 2020; Dahbura et al., 2021; Martínez Dahbura et al., 2023). We note that these methods were first developed for undirected graphs, and their application to directed networks necessitates non-trivial extensions of the computational machinery, providing an additional contribution of our work. [^7] Our methods are implemented in the scalable, open-source, and platform-independent R package bigergm, which is available on CRAN servers (R Core Team, 2024). Implementation details can be found in the Appendix.
@@
Our main result is that coders exert a negative externality on other coders when creating a dependency on another package. Other coders are then also less likely to create a dependency with that package. As a result, the network is relatively sparse. The interdependence among libraries implies that there exists a risk that a vulnerability in a single package has large adverse consequences for the entire ecosystem. One example of how a vulnerability in one package impacted a significant part of critical web infrastructure is Heartbleed, the infamous bug in the widely used SSL/TLS cryptography library, resulting from an improper input variable validation (Durumeric et al., 2014). When Heartbleed was disclosed in 2014, it rendered up to 55% of secure webservers vulnerable to data theft, including of the Canadian Revenue Agency, and Community Health Systems, a large U.S. hospital chain. Our model provides a further argument for ensuring the security of highly interdependent software packages. Not only can a vulnerability in such a package affect a relatively large fraction of the entire ecosystem, because of the externality we identify, it is also likely that this fraction increases as time goes on.
@@
Another important question in the study of network formation processes is whether linked packages are similar or dissimilar with respect to their covariates. We find evidence of similarity among the same type of software packages in terms of their Size, Popularityand Maturityas well as between different-type software packages in terms of their Popularity. The effect of similarity decreases with the (log of) the number of packages of the same unobservable type. We further find evidence of dis-similarity between different-type software packages in terms of their Maturity  and Size. In other words, mature, popular and large software packages of one type are likely to depend on similar packages of the same type, but on less mature and smaller software packages of another type. This is intuitive, because mature, popular, and large packages are likely to have a large user base with high expectations of the software and coders try to satisfy this demand by providing sophisticated functionality with the help of mature, popular, and large software packages of the same type. These packages are often built using a lot of low-level functionality from large but fairly generic libraries, which is why large popular and mature software packages of one type depend on larger, but less popular and less mature software packages of another type. One example of this is the inclusion of a payment gateway in an e-commerce application. The e-commerce application itself can be sophisticated and complex, like ebay is in the provision of their auction mechanism. For such an application it is particularly valuable to provide users with the ability to use a variety of different payment methods, including credit card, paypal, or buy now pay later solutions like Klarna. Our result also aligns with the recent trend in software development away from large monolithic applications and towards interconnected microservices (Traore et al., 2022).
Lastly, we adapt the susceptible-infected-recovered (SIR) model from epidemiology to examine the spread of vulnerabilities in the Python Pypi dependency graph. We measure a repository’s $k$-step systemicness as the number of downstream packages that are potentially rendered vulnerable by a vulnerability or bug in the upstream repository.
We then study the efficacy of targeted interventions to mitigate vulnerability spread, drawing parallels with vaccination strategies in public health. By focusing on securing the most critical nodes—determined according to in-degree, expected fatality or a combination of betweenness centrality and expected fatality—we assess the potential to significantly reduce the risk of vulnerability contagion.
Our findings suggest that protecting even ten percent of the most critical repositories may reduce the average systemic risk, but cannot completely flatten out the risk profile of the ecosystem. On the other hand, we argue that the introduction of AI-assisted coding may have a larger effect on reducing systemic risk. Recent work has shown that AI tools such as GitHub Copilot can increase productivity of software developers (Peng et al., 2023; Edelman et al., 2023; Noy and Zhang, 2023). In our model, this corresponds to a decrease of the cost of producing software in-house, without relying on creating dependencies or creating fewer dependencies. This can be modeled as an increase of the relative cost of a dependency. By simulating different magnitudes of the productivity improvement from AI-assisted coding, we show that the average systemic risk profile is flattened out by an increase of cost of merely five percent. The mechanism through which this risk reduction is achieved is the brute-force decrease in density for the network implied by increases in costs of dependencies. Overall, our results help better understand the various driving forces and motifs in software development and how they shape the network of software dependencies.
@@
The package manager model is nowadays adopted by most programming languages which makes it feasible to use dependency graphs to study a wide variety of settings. Kikas et al. (2017), for example, study the structure and evolution of the dependency graph of JavaScript, Ruby, and Rust. The authors emphasize that dependency graphs of all three programming languages become increasingly vulnerable to the removal of a single software package. An active literature studies the network structure of dependency graphs (Decan et al., 2019) to assess the vulnerability of a software ecosystem (see, for example, Zimmermann et al., 2019).
@@
These papers show the breadth of literature studying open source ecosystems and dependency graphs. However, the literature considers the network either as stochastic or even as static and given. In contrast, we model the formation of dependencies as coders’ strategic choice in the presence of various and competing mechanisms that either increase or reduce utility. [^10] Blume et al. (2013) study how possibly contagious links affect network formation in a general setting. While we do not study the consequences of the externality we identify for contagion, this is a most worthwhile avenue for future research. Perhaps the closest work is Boysel (2023), that uses data from a sample of repositories to estimate a structural model of network formation and maintainers effort. In his model, maintainers sequentially update their dependencies, then play a network game where their efforts best responses are a function of other maintainers efforts. While related, our approaches are different, since Boysel (2023) relies on sample panel data and the analysis focuses on the dynamics of myopic best-response. In our paper we focus on the long-run equilibrium and the cross-section of the dependencies among all repositories at a particular point in time.
@@
The theoretical literature on strategic network formation has pointed out the role of externalities and heterogeneity in shaping the equilibrium networks (Bramoullé et al., 2016; Galeotti et al., 2006; Jackson and Wolinsky, 1996). Estimating strategic models of network formation is a challenging econometric task, because the presence of externalities implies strong correlations among links and multiple equilibria (Mele, 2017; Snijders, 2002; DePaula et al., 2018; Chandrasekhar, 2016; DePaula, 2017; Boucher and Mourifie, 2017; Graham, 2017, 2020). In this paper, we model network formation as a sequential process and focus on the long-run stationary equilibrium of the model (Mele, 2017, 2022). Because the sequential network formation works as an equilibrium selection mechanism, we are able to alleviate the problems arising from multiple equilibria.
@@
Adding unobserved heterogeneity further complicates identification, estimation and inference (Schweinberger and Handcock, 2015; Graham, 2017; Mele, 2022). Other works have considered conditionally independent links without externalities (Graham, 2017, 2020; DePaula, 2017; Chandrasekhar, 2016), providing a framework for estimation and identification in random and fixed effects approaches. On the other hand, because link externalities are an important feature in this context, we move away from conditionally independent links, and model nodes’ unobserved heterogeneity as discrete types, whose realization is independent of observable characteristics and network, in a random effect approach. We can thus adapt methods of community discovery for random graphs to estimate the types, extending the work of Babkin et al. (2020) and in Dahbura et al. (2021); Martínez Dahbura et al. (2023) to accommodate for directed networks. Our two-steps method scales well to large networks, thus improving the computational challenges arising in estimation of these complex models (Boucher and Mourifie, 2017; Vu et al., 2013; Bonhomme et al., 2019).
@@
Our model is able to estimate the magnitude of externalities as well as homophily (Currarini et al., 2010; Jackson, 2008; Chandrasekhar, 2016), the tendency of individuals to form links to similar individuals. Furthermore, our model is able to detect heterophily (or competition) among maintainers. More specifically, we allow the homophily to vary by unobservables, while in most models the homophily is estimated only for observable characteristics (Currarini et al., 2010; Schweinberger and Handcock, 2015; DePaula et al., 2018; Chandrasekhar, 2016; Graham, 2020).